Accelerator Update for January 30, 2008 - February 1, 2008
Wednesday January 30
The day shift began with Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (9.8E10), with the Antiproton Source stacking (163.4E10), with the Main Injector (MI), Linac, and Booster being prepared for access, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments off due to accesses.
Operations established store 5873 at 8:01 AM with an initial luminosity of 229.6E30.
Operations reported at 8:19 AM that Linac and NuMI had already begun their accesses, while Booster and MI were still being prepared for access. This 10 – 12 hour shutdown was being driven by MI’s need to replace a quadrupole magnet (Q631).
After the Linac work was completed, Operations established beam to the Linac absorber at 12:13 PM. Linac experts began a LRF5 study that lasted until the end of the day shift.
At 12:23 PM, a Linac Klystron RF station (KRF6) tripped off. It reset.
Booster came out of access at 12:49 PM.
Operations status report at 1:25 PM: the MI Q631 magnet was disconnected and removed from the tunnel, the MI LCW leak on Q101 was repaired, vacuum experts were checking for leaks on the MI RF sections, and the shielding for the 180 degree magnet was removed so experts could inspect the NMR probe.
At 3:41 PM, KRF7 tripped off. It reset.
NuMI came out of access at 4:24 PM.
At 5:06 PM, KRF6 tripped off. It reset.
At 8:33 PM, KRF7 tripped off. It reset.
At 8:39 PM, a TeV RF station (TRF5) tripped off on a power amplifier (PA) LCW off indication, but there was no problem with the water pumps. The station will remain off until the store ends.
Booster studies ended at 9:28 PM.
A MI expert reported at 11:11 PM that the vacuum in MI-10 was good enough to send beam.
Thursday January 31
The midnight shift began with monitoring store 5873 (59.98E30), with Pbar stacking (159.5E10), and with the MI turn on in progress.
Operations reported at 1 AM that NuMI, Pbar, and MI were ready for beam.
At 1:02 AM, MI suffered a vacuum burst. It reset.
At 1:05 AM, KRF4 tripped off. It reset.
At 1:24 AM, Operations reported a problem with a BNB power supply (E:HV860) that had an output that didn’t agree with its setting. An expert was called in to investigate.
At 1:25 AM, Operations reported that the MI interlock Bypass Loop had tripped due to a problem at MI-10. The Loop would trip off and then reset. An expert investigated.
Operations established beam to BNB at 2 AM.
At 5:05 AM, KRF7 tripped off. It reset.
At 5:14 AM, LRF3 tripped off and would not reset. Operators called in a Linac expert who repaired the station.
Pbar began stacking and NuMI was taking beam by 6:48 AM.
At 7:22 AM, Operations reported that stack was lost when an Accumulator shunt (A:BS103) went to zero. A Pbar expert investigated and reported that the problem was in the tunnel. Operators prepared the area for an access.
Operations reported at 8 AM that the Recycler had stashed antiprotons once during the midnight shift.
A Pbar expert fired the Titanium Sublimation Pump (TSP) in the Accumulator at 8:15 AM.
At 8:20 AM, BRF10 tripped off. It reset.
Pbar experts came out of access at 8:26 AM.
Switchyard (SY) experts began using beam for studies at 9:09 AM.
At 9:34 AM, LRF5 tripped off. Operators investigated and found its breaker tripped. Operators contacted Linac experts.
Linac resumed sending beam at 9:51 AM.
At 11:40 AM, a Booster vacuum pump tripped off (B:IPRF17) and took down BRF17 & 18. Everything reset.
At 11:45 AM, KRF3 tripped off. It reset.
An Instrumentation expert rebooted the VME for the Recycler and MI DCCTs at 1:22 PM.
Operations terminated store 5873 at 2 PM.
A TeV RF expert began work on TRF5 at 2:59 PM.
Operations put the TeV into shot setup at 3:01 PM.
At 3:22 PM, BRF10 tripped off. It reset.
Operations reported at 4 PM that the Recycler had stashed once during the day shift.
Operations established store 5874 at 5:08 PM with an initial luminosity of 151.69E30.
SY experts stopped using beam at 5:21 PM.
At 6:26 PM, LRF3 had another problem with its gradient staying at zero, but fortunately there was a diagnostic scope hooked up to the station that recorded some noise.
At 7:12 PM, KRF7 tripped off. It reset.
At 7:14 PM, TRF3 tripped off.
At 8:09 PM, BRF10 tripped off for the first of two times in about 50 minutes. It reset.
Operations reported at midnight that the Recycler had stashed twice during the evening shift.
Friday February 1
The midnight shift began with Operations monitoring the store 5874 (63.39E30) and stash (127.1E10), with Pbar stacking (32.5E10), and with NuMI and BNB taking beam.
At 2:26 AM, KRF5 tripped off. It reset.
At 4:02 AM, KRF3 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today are to end store 5874 around 10 AM and then allow CDF and D0 to make an access. The Run Coordinator reported that January was the second best month for luminosity with a total of 164.9 pb-1 (inverse picobarns). The record is 166.8 pb -1 for one month.
Comments and Suggestions
|last modified 02/01/2008 email Fermilab|
|Security, Privacy, Legal|